# Training–free AI for Earth Observation Change Detection using Physics Aware Neuromorphic Networks

### v1

@@ Methods
The images are level 1C pre–processed multi-spectral images from the Multi-Spectral Imager (MSI) camera on-board the Sentinel-2 Earth observation satellite constellation. Only the 10 highest spatial resolution bands are used in this study. Table 3 shows the specific bands used in this study along with the corresponding index number used to reference the bands. The dataset covers four categories of natural disasters: fires, floods, hurricanes and landslides. Each category includes five natural disaster events, except for floods which have four, giving a total of 19 natural disaster events. For each natural disaster event there are five sequential images; the first four are all before the natural disaster event and the final image is after. Along with each natural disaster event, there is an accompanying change mask that contains three labels: unaffected regions, affected areas, and clouds from the fourth and fifth images. An example of a hurricane disaster event is shown in Fig. 4, where damage to the island vegetation can be seen in the fifth image and the change mask image shows the unaffected regions in purple, the affected region in aqua and the clouds in yellow. The change mask in this study is only used to evaluate the performance of the proposed model.
Band B2 B3 B4 B5 B6 B7 B8 B8a B11 B12
Index (m) 01 02 03 04 05 06 07 08 09 10
Table 3: Sentinel-2 bands used in this study and the corresponding index used to reference each band.
The data was preprocesessed by first taking the log of the rgb pixel values and then scaling the values to the interval [-1, +1] for each band in the data. This process is described by equation (1):
@@
where the min and max values for each band were chosen to be consistent with Ref. [21], to allow a direct comparison with their state-of-the-art variational autoencoder model. Missing values were set to a value just above zero of 0.005.
Figure 4: An example of a hurricane natural disaster scene, with the first 4 images before the hurricane and image 5 is after. The change mask (last column) shows the affected (aqua) and unaffected (purple) regions along with the clouds (yellow) in the images directly before and after the hurricane.
@@
The PANN is based on a physically–motivated model of a physical neuromorphic network comprised of self–organised nanowire networks [41, 43, 45]. This is done by first creating a nanowire network and then converting it into a graph representation which is used in the PANN. This process is shown in Fig. 5. In this study, the network for the PANN was created by distributing 803 nanowires over a $158\times 158\,\mu$m 2D plane with nanowire centres sampled from a generalised normal distribution with a beta value of 5 and with nanowire orientations sampled on $[0,\pi]$. Nanowire lengths were sampled from a Gaussian distribution with an average of $30\,\mu$m and a standard deviation of $6\,\mu$m. As nanowire networks are nano-electronic systems, input signals are delivered via electrodes, which were modelled with a diameter $5\,\mu$m and with evenly spaced placements in a $16\times 16$ grid over the 2D plane with a margin of $15\,\mu$m, giving a pitch of $8\,\mu$m between electrodes. Where the nanowires overlap with other nanowires or electrodes, electrical junctions are formed whose internal states evolve in time, as described below. Figure 5(a) shows the physical nanowire network, with the nanowires displayed as grey lines and nanowire–nanowire junctions as small grey dots. The electrodes are shown as the large green circles and the nanowire–electrode junctions are shown as small dark green dots. Figure 5(b) shows the corresponding graphical representation of the network used in the PANN. The black nodes correspond to the nanowires and the green nodes correspond to the electrodes, used as input nodes. The edges correspond to the electrical junctions. The resulting network had a total of 1059 nodes and 12,279 edges. 256 nodes were selected as input nodes and from the remaining 803 nodes, 400 were selected at random as output nodes (note that in a real physical device, electrodes would also serve as output nodes).
Figure 5: (a) PANN model nanowire representation. The nanowires are the grey lines and the electrodes are the large green circles. Nanowire–nanowire junctions are shown as grey dots and nanowire–electrode junctions are shown as dark green dots. (b) graph representation of the PANN. Green nodes are the input nodes which correspond to the electrodes. Black nodes correspond to the nanowires and edges correspond to the junctions, nanowire–nanowire and nanowire–electrode junctions.
@@
with all edges set to an initial state $\lambda(t=0)=0$. In this study, the following parameters were used: $V_{\text{reset}}=5\times 10^{-3}\,$, $V_{\text{set}}=1\times 10^{-2}\,$ and $\lambda_{\text{max}}=1.5\times 10^{-2}\,$ s. Full details of the model and its validation against experimental nanowire networks are provided in Refs. [43, 45]. In contrast to artificial neural networks, the dynamic nature of these physical neuromorphic networks means that weights are not trained, which not only makes them more energy efficient, but also allows all the data to be used for evaluating the model.
@@
Figure 6: Diagram of the PANN workflow showing the pipeline from the input images to the output features, $F^{a,b}(t)$ from the networks. First, automatic feature engineering is applied to the input images to determine which bands of the images to use. The images are then broken into a sequence of small $32\times 32$ pixel tiles, $x^{a,b}_{N}(t)$, where $a,b$ is the location of the tile and $N$ is the number of bands. The creation of the tile sequence is shown by the coloured (red, orange, yellow, lime and green) boxes and are then fed into the max pooling layer. The bands are separated, as denoted by $\bigotimes$, with each band being fed into a unique PANN model. In this example PANN01-03 and PANN09 are used. The outputs from all the PANNs that are used are concatenated together, as denoted by $\bigoplus$, to produce the feature sequence $F^{a,b}(t)$ for tile $x^{a,b}_{N}(t)$.
To determine whether a change from a natural disaster occurred at a given tile location, the distance between the images of the tile sequence is measured in the feature space. It is assumed that if there is any change it is in the last image of the sequence, so only the distances between the final image and each of the previous images is compared. This is the same change score method as introduced in Ref. [21]. Formally, the change score, $S$, is given by
@@
where $dist$ is an arbitrary distance metric. In this study we focus on three different distance metrics: Euclidean, Cosine and Correlation. Unless otherwise stated, the Correlation distance was used. The Cosine distance is an angle-base metric and is given by
$\text{CosineDist}(\mathbf{u},\mathbf{v})=1-\frac{\mathbf{u}\cdot\mathbf{v}}{||\mathbf{u}||*||\mathbf{v}||}\qquad.$ (4)
@@
The automatic feature engineering process is performed once for each natural disaster scene and is performed first, before the normalisation and tiling stages. The feature engineering process leverages common indices such as the normalised burn ratio index [57] and the normalised difference vegetation index [58] to calculate natural disaster class scores. Using these scores, a binary decision tree is followed, as shown in Fig. 7, to determine what class of natural disaster is in the scene. The score for each natural disaster class is calculated by creating an index image for the images directly before and after the natural disaster event using equation (6), where $BX$ and $BY$ are predetermined bands depending on what type of natural disaster is being assessed. The images are converted to a binary image using a predefined threshold value and a difference image is created by a pixel–wise comparison. The final score is the percentage of pixels that record a change. Once a natural disaster type has been selected, the feature engineering process then selects a subset of bands to be used in the rest of the model depending on the natural disaster class. The specific values and bands used for the natural disaster class score are shown in Fig. 7.
@@
We are releasing the full code alongside this paper at https://github.com/ssmi9157/ChangeDetectionPANN. The dataset analysed in this study is available in the RaVAEn repository, https://github.com/spaceml-org/RaVAEn.
@@ Acknowledgements
Appendix
Figure 8: The 2D UMAP projection of the pixel space for a flood natural disaster event: Left – position of the before (blue) and after (orange) tiles and with percentage change in the after tiles indicated by the colour bar; Right – the corresponding tile images.
Figure 9: Visual comparison between the 2D UMAP projection of the PANN feature space for a hurricane natural disaster, for the full tile sequence, using three different distance metrics: Euclidean, Cosine and Correlation distances.

### v2

@@ Methods
The primary dataset used in this study is a time series of multi–spectral EO of natural disasters released by Růžička et al. [28]. The images are level 1C pre–processed multi-spectral images from the Multi-Spectral Imager [67] camera on-board the Sentinel-2 Earth observation satellite constellation. Only the 10 highest spatial resolution bands are in the dataset. All bands have a spatial resolution of 10 m. The bands that are not naturally at 10 m resolution were interpolated to the higher resolution. Table 4 shows the specific bands used in this study along with the corresponding index number used to reference the bands. The dataset covers four categories of natural disasters: fires, floods, hurricanes and landslides. Each category includes five natural disaster events, except for floods which have four, giving a total of 19 natural disaster events. The individual samples were created by tiling each event, with no overlap, into $32\times 32$ pixel sequence tiles, resulting in a total of 62,655 samples used to evaluate the PANN model. For each natural disaster event there are five sequential images; the first four are all before the natural disaster event and the final image is after. Along with each natural disaster event, there is an accompanying change mask that contains three labels: unaffected regions, affected regions, and clouds from the fourth and fifth images. An example of a hurricane disaster event is shown in Fig. 5, where damage to the island vegetation can be seen in the fifth image and the change mask image shows the unaffected regions in purple, the affected region in aqua and the clouds in yellow. The change mask in this study is only used to evaluate the performance of the proposed model.
In this study, we created a second dataset to evaluate the performance of the PANN from a different sensor, the LandSat-8 Operational Land Imager [68]. This was done by selecting the natural disaster events in the Sentinel-2 dataset that did not have any clouds obscuring the change regions in the change mask. This left 8 natural disaster events: 2 fires, 2 floods and 4 landslides. The Top of Atmosphere images from the LandSat-8 satellite with the 9 highest spatial resolution bands of 30 m were used. To create the time series of images for each natural disaster event, only images with 40% cloud cover or less were used. Like the Sentinel-2 dataset, each natural disaster event has five sequential images, with the first four before the event and the final image is after. The corresponding change mask from the Sentinel-2 dataset was used after resizing to 30 m resolution using the nearest-neighbour method and adding the clouds in the fourth and fifth images only. Clouds were detected automatically using the Google Earth Engine LandSat simple cloud score algorithm [69], with a score $\geq 55$% considered a cloud. As with the Sentinel-2 dataset, individual samples were created by tiling each event, with no overlap, into $16\times 16$ pixel sequence tiles, resulting in a total of 8,199 samples.
Satellite Index (m)
01 02 03 04 05 06 07 08 09 10
Sentinel-2 B2 B3 B4 B5 B6 B7 B8 B8a B11 B12
LandSat-8 B2 B3 B4 B5 B6 B7 B9 B10 B11
Table 4: Sentinel-2 and LandSat-8 bands used in this study and the corresponding index used to reference each band.
The data was preprocessed by first taking the log of the RGB pixel values and then scaling the values to the interval [-1, +1] for each band in the data. This process is described by equation (1):
@@
where the min and max values for each band are fixed and were chosen to be consistent with Ref. [28] for the Sentinel-2 dataset, to allow a direct comparison with their state-of-the-art variational autoencoder model. A different set of min and max values were manually selected by visual inspection for the LandSat-8 dataset. Missing values were set to a value just above zero of 0.005.
Figure 5: An example of a hurricane natural disaster scene, with the first 4 images before the hurricane and image 5 is after. The change mask (last column) shows the affected (aqua) and unaffected (purple) regions along with the clouds (yellow) in the images directly before and after the hurricane. Satellite images are from Sentinel-2 complied by Růžička et al. and available in their git[57].
@@
[…]
Figure 6: (a) PANN model nanowire representation. The nanowires are the grey lines and the electrodes are the large green circles. Nanowire–nanowire junctions are shown as grey dots and nanowire–electrode junctions are shown as dark green dots. (b) graph representation of the PANN. Green nodes are the input nodes which correspond to the electrodes. Black nodes correspond to the nanowires, while edges correspond to the nanowire–nanowire and nanowire–electrode junctions.
@@
with all edges set to an initial state $\lambda(t=0)=0$. In this study, the following parameters were used: $_{\text{reset}}=5\times 10^{-3}\,V, V$$_{\text{set}}=1\times 10^{-2}\,V$ and $\lambda_{\text{max}}=1.5\times 10^{-2}\,$V s. Full details of the model and its validation against experimental nanowire networks are provided in Refs. [51, 53]. In contrast to artificial neural networks, the dynamic nature of these physical neuromorphic networks means that the weights are not trained, which not only makes them more energy efficient, but also allows all the data to be used for evaluating the model.
@@
Figure 7: Diagram of the PANN workflow showing the pipeline from the input images to the output features, $F^{a,b}(t)$ from the networks. First, automatic feature engineering is applied to the input images to determine which bands of the images to use. The images are then broken into a sequence of small $32\times 32$ pixel tiles, $x^{a,b}_{N}(t)$, where $a,b$ is the location of the tile and $N$ is the number of bands. The creation of the tile sequence is shown by the coloured (red, orange, yellow, lime and green) boxes and are then fed into the max pooling layer. The bands are separated, as denoted by $\bigotimes$, with each band being fed into a unique PANN model. In this example PANN01-03 and PANN09 are used. The outputs from all the PANNs that are used are concatenated together, as denoted by $\bigoplus$, to produce the feature sequence $F^{a,b}(t)$ for tile $x^{a,b}_{N}(t)$. Satellite images are from Sentinel-2 complied by Růžička et al. and available in their git[57].
To determine whether a change from a natural disaster occurred at a given tile location, the distance between the images of the tile sequence is measured in the feature space. It is assumed that if there is any change it is in the last image of the sequence, so only the distances between the final image and each of the previous images are compared. This is the same change score method as introduced in Ref. [28]. Formally, the change score, $S$, is given by
@@
where distdist is an arbitrary distance metric. In this study we focus on three different distance metrics: Euclidean, Cosine and Correlation. Unless otherwise stated, the Correlation distance was used. The Cosine distance is an angle-base metric and is given by
[…]
@@
The automatic feature engineering process is performed once for each natural disaster scene and is performed first, before the normalisation and tiling stages. The feature engineering process leverages common indices such as the normalised burn ratio index [72] and the normalised difference vegetation index [73] to calculate natural disaster class scores. Using these scores, a binary decision tree is followed, as shown in Fig. 8, to determine what class of natural disaster is in the scene. The score for each natural disaster class is calculated by creating an index image for the images directly before and after the natural disaster event using equation (6), where $BX$ and $BY$ are predetermined bands depending on what type of natural disaster is being assessed. The images are converted to a binary image using a predefined threshold value and a difference image is created by a pixel–wise comparison. The final score is the percentage of pixels that record a change. Once a natural disaster type has been selected, the feature engineering process then selects a subset of bands to be used in the rest of the model depending on the natural disaster class. The specific values and bands used for the natural disaster class score are shown in Fig. 8. See Supplementary Fig. S4 online for the specific values and bands used for the LandSat-8 data.
@@
We are releasing the full code alongside this paper at https://github.com/ssmi9157/ChangeDetectionPANN. The Sentinel-2 dataset analysed in this study is available in the RaVAEn repository https://github.com/spaceml-org/RaVAEn and the LandSat-8 dataset is available at https://zenodo.org/records/16916355.